### v2

# Certifiable Deep Learning for Reachability Using a New Lipschitz Continuous Value Function

Abstract
We propose a new reachability learning framework for high-dimensional nonlinear systems, focusing on reach-avoid problems. These problems require computing the reach-avoid set, which ensures that all its elements can safely reach a target set despite any disturbance within pre-specified bounds. Our framework has two main parts: offline learning of a newly designed reach-avoid value function and post-learning certification. Compared to prior works, our new value function is Lipschitz continuous and its associated Bellman operator is a contraction mapping, both of which improve the learning performance. To ensure deterministic guarantees of our learned reach-avoid set, we introduce two efficient post-learning certification methods. Both methods can be used online for real-time local certification or offline for comprehensive certification. We validate our framework in a 12-dimensional crazyflie drone racing hardware experiment and a simulated 10-dimensional highway takeover example.
Index Terms:
Reachability, nonlinear control, deep reinforcement learning.
I Introduction
Ensuring the safe and reliable operation of robotic systems in uncertain environments is a critical challenge. For instance, humanoid robots must safely navigate around humans while performing their tasks, and air taxi vehicles need real-time synthesis of safe, collision-free trajectories in crowded airspace. In these safety-critical applications, it is essential to complete various tasks safely despite unpredictable disturbances. Additionally, knowing the set of states from which the task can be safely completed is important because the task can fail once the robot exits that set. Reachability analysis addresses this challenge by computing the optimal control and determining the reach-avoid set—a set of states that can safely reach a target set under all possible disturbances within a specified bound.
Traditional Hamilton-Jacobi reachability analysis methods, such as [2, 3, 4], leverage dynamic programming to synthesize the optimal control and a reachability value function, whose sign indicates whether or not a state can safely reach the target set. Though theoretically sound, they suffer from the curse of dimensionality [5]. As the system’s dimension increases, the computational complexity grows exponentially, making these methods impractical for real-world applications without approximation or further logic to manage the problem size.
There has been interest in leveraging machine learning techniques to learn reachability value functions for high-dimensional systems [6, 7, 8, 9, 10, 11, 12]. However, a major drawback of existing learning-based approaches is the lack of deterministic guarantees. Without such guarantees, it is difficult to ensure the reliability and safety of the robotic systems in critical scenarios.
Recent works [13, 14] provide probabilistic safety guarantees for the learned reach-avoid sets. Additionally, safety filter approaches [11, 12, 15] have been proposed, which offer point-wise guarantees by ensuring safety for individual states.
In this work, we asked if it is possible to learn a set of states that have deterministic reach-avoid guarantees. We propose a new framework for learning trustworthy reach-avoid sets for high-dimensional systems. Our framework involves first learning a newly designed reach-avoid value function, and then conducting set-based certification to ensure that all states in a certified reach-avoid set safely reach the target set despite any disturbance. Specifically:
Figure 1: Applying our reachability analysis framework to drone racing. In (a), hardware experiments demonstrate that our learned control policy enables an ego drone to safely overtake another drone, despite unpredictable disturbances in the other drone’s acceleration. In (b), we illustrate the concept of the propeller induced airflow [16], which can affect other drones’ flight. In (c), we apply our learned control policy in a simulation with randomly sampled disturbances. In (d), we project the learned reach-avoid value function onto the $(x,y)$ position of the ego drone. The super-zero level set, outlined by dashed curves, indicates our learned reach-avoid (RA) set. In (e), we plot the certified RA sets using Lipschitz and second-order cone programming certification.
  1. 1.
    We propose a new reach-avoid value function that is provably Lipschitz continuous, and its Bellman operator is a contraction mapping, removing the need for contractive Bellman operator approximation commonly used in prior works. Moreover, we demonstrate that the optimal control derived from our value function tends to reach the target set more rapidly than those from prior works. We apply deep reinforcement learning (RL) to learn this new value function for high-dimensional nonlinear systems.
  2. 2.
    We develop two reach-avoid set certification methods. The first uses the Lipschitz constant of the dynamics to certify the safety of a subset of the learned reach-avoid set, ensuring all its elements can safely reach the target set amid potential disturbances. The second employs second-order cone programming to do the same. Both methods offer deterministic assurances. They can be applied online to verify if a neighboring set around the current state can safely reach the target set despite any disturbance within a pre-specified set, or offline for comprehensive certification.
  3. 3.
    We demonstrate the computational benefits of our new value function and the assurance of our (real-time) certification methods through hardware experiments and simulations. We show empirically that our learned policy compares favorably with that of prior works, and that the Lipschitz continuity of our new value function can accelerate value function learning.
II Related works
Reachability learning methods. DeepReach [6] is a pioneering work on learning finite-horizon reachability value functions. In other studies, such as [7, 8, 10, 9, 11, 15, 12], the assumption of a known horizon is relaxed and infinite horizon reachability learning problems are considered. In this work, we also consider the infinite horizon case. We introduce a new value function which is provably Lipschitz continuous and whose Bellman operator is a contraction mapping, offering favorable computational efficiency when compared with prior work.
Verification of learning-based control. Recent work [13, 14, 17] has provided probabilistic safety guarantees for DeepReach; in this paper, we introduce methods that provide deterministic reach-avoid guarantees and we show how they could be used locally in real time. Other studies [11, 15, 12] provide point-wise safety filters. However, we propose set-based reach-avoid certification methods to verify if all states in a set can safely reach the target set amid potential disturbances. Our certification methods also differ from existing set-based approaches for verifying neural network-controlled systems, including those that verify regions of attraction [18, 19, 20], forward reachability sets [21, 22, 23, 24], and safe sets using barrier certificates [25, 26, 27]. In general, these methods do not certify if a set of states is within the ground truth reach-avoid set, as ours do.
Constrained optimal control. Control barrier functions (CBFs) offer safety guarantees [28, 29, 30], but they tend to be more conservative than model predictive control (MPC) [31, 32], which optimally balances task performance and safety. However, constructing or learning CBFs and solving MPC can be difficult for nonlinear systems with complex constraints. By leveraging deep neural networks, constrained reinforcement learning (CRL) [33, 34, 35] learns control policies that maximize task rewards while adhering to complex constraints. Nonetheless, the positive sign of a CBF provides safety information for a state; it does not indicate whether a trajectory from that state can reach the target set. Similarly, the value functions in MPC and CRL do not provide information about whether a state can safely reach the target set. In contrast, our new value function not only provides the optimal control for the worst-case disturbance, but also indicates, based on its sign, whether or not a state is in the reach-avoid set.
III Problem Formulation
We consider the uncertain nonlinear dynamics described by
$x_{t+1}=f(x_{t},u_{t},d_{t}),$ (1)
where $x_{t}\in\mathbb{R}^{n}$ is the state, $u_{t}\in{\mathcal{U}}\subseteq{\mathbb{R}^{m_{u}}}$ is the control, and $d_{t}\in{\mathcal{D}}\subseteq{\mathbb{R}^{m_{d}}}$ represents the disturbance, such as model mismatches or uncertain actions of other agents. We assume both ${\mathcal{U}}$ and ${\mathcal{D}}$ are compact and continuous sets. We define a state trajectory originating from $x_{0}$ under a control policy $\pi(\cdot):\mathbb{R}^{n}\to{\mathcal{U}}$ and a disturbance policy $\phi(\cdot,\cdot):\mathbb{R}^{n}\times{\mathcal{U}}\to{\mathcal{D}}$ as ${\xi_{x_{0}}^{\pi,\phi}}:=\{x_{t}\}_{t=0}^{\infty}$, where $x_{t+1}=f(x_{t},{\pi}(x_{t}),{\phi}(x_{t},{\pi}(x_{t})))$$\forall t$.
Let ${\mathcal{T}}\subseteq\mathbb{R}^{n}$ be an open set, representing a desired target set. We assume that there exists a Lipschitz continuous, bounded reward function $r(x)$ indicating if a state $x$ is in the target set,
$r(x)>0\Longleftrightarrow x\in{\mathcal{T}}.$ (2)
We consider a finite number of Lipschitz continuous, bounded constraint functions $c_{i}(x)>0,\forall i\in{\mathcal{I}}$. We can simplify the representation of constraints by considering their pointwise minimum $c(x):=\min_{i\in{\mathcal{I}}}c_{i}(x)$. We define the constraint set as ${\mathcal{C}}:=\{x:c(x)>0\}$, and we have
$c(x)>0\Longleftrightarrow x\in{\mathcal{C}}.$ (3)
We look for states that can be controlled to the target set safely under the worst-case disturbance, with dynamics given by (1). We refer to this set as the reach-avoid (RA) set [36, 3]:
:=${\mathcal{R}}:=\mathopen{}\mathclose{{}\left\{\begin{aligned} x_{0}:&\exists\pi\ \textrm{such that }\forall\phi,\ \exists T<\infty,\\ &\big{(}r(x_{T})>0\ \wedge\ \forall t\in[0,T],c(x_{t})>0\big{)}\end{aligned}}\right\},$ (4)
which includes all the states that can reach the target set safely in finite time despite any disturbance within the set ${\mathcal{D}}$.
In this paper, we compute the RA set for high-dimensional nonlinear systems with complex constraints.
Running example, safe take-over in drone racing: We model the drone take-over example in Figure 1 as an RA problem, where two crazyflie drones [37] compete to fly through an orange gate. The first drone (ego agent) starts behind and aims to overtake the other drone. The second drone flies directly to the gate using an LQR controller, but its acceleration is uncertain to the first drone. We model the uncertain part of the other drone’s acceleration by a disturbance $\|d_{t}\|_{2}\leq\varepsilon_{d}:=0.1\textrm{ m/s}^{2}$. We compute the RA set to ensure the ego drone can safely overtake the other despite this disturbance.
We consider a 12-dimensional dynamics [38], where the $i$-th drone’s state is $x_{t}^{i}=[p_{x,t}^{i},v_{x,t}^{i},p_{y,t}^{i},v_{y,t}^{i},p_{z,t}^{i},v_{z,t}^{i}]$. In each of the $(x,y,z)$ axes, the $i$-th drone is modeled by double integrator dynamics, and the control is its acceleration $u_{t}^{i}=[a_{x,t}^{i},a_{y,t}^{i},a_{z,t}^{i}]$, with $\|u_{t}^{i}\|_{\infty}\leq\varepsilon_{u}:=1\textrm{m/s}^{2}$. We model the center of the gate as the origin. The radius of the orange gate is 0.3 meters, and the radius of the crazyflie drone is 0.05 meters. We consider a target set for the ego drone, as plotted in Figure 1:
${\mathcal{T}}=\mathopen{}\mathclose{{}\left\{\begin{aligned} x:\ &p_{y}^{1}-p_{y}^{2}>0,\hskip 14.22636ptv_{y}^{1}-v_{y}^{2}>0,\\ &|p_{x}^{1}|<0.3,\hskip 21.33955pt|p_{z}^{1}|<0.3\end{aligned}}\right\}.$ (5)
To ensure the ego drone flies through the gate, we constrain:
±$\pm p_{x,t}^{1}-p_{y,t}^{1}>-0.05,\hskip 14.22636pt\pm p_{z,t}^{1}-p_{y,t}^{1}>-0.05.$ (6)
To ensure safe flight, the ego drone should avoid the area affected by the airflow from the other drone, as depicted in Figure 1, using the constraint:
$\mathopen{}\mathclose{{}\left\|\begin{bmatrix}p_{x,t}^{1}-p_{x,t}^{2}\\ p_{y,t}^{1}-p_{y,t}^{2}\end{bmatrix}}\right\|_{2}^{2}>\Big{(}1+\max(p_{z,t}^{2}-p_{z,t}^{1},0)\Big{)}\times 0.2,$ (7)
where the required separation distance between the ego drone and the other drone increases as their height difference grows.
Numerically computing the RA set directly for this problem is computationally infeasible [39]. We introduce our new reachability learning method in the following sections.
IV A New RA Value Function
In this section, we propose a new RA value function for evaluating if a state belongs to the RA set. Unlike prior works [7, 8, 10, 9, 11, 15, 12], our value function incorporates a time-discount factor. This results in a Lipschitz-continuous value function, which appears to accelerate the learning process, and establishes a contractive Bellman equation, eliminating the need for the contractive Bellman equation approximation commonly used in prior works. Furthermore, we show that the optimal control derived from this new value function tends to reach the target set rapidly.
We begin the construction of our new value function by first introducing the concept of RA measure, which assesses whether a trajectory can reach the target set safely. Let ${\xi_{x_{0}}^{\pi,\phi}}$ be a trajectory that enters the target set safely at a stage $t$. We have ${r}({x_{t}})>0$ and ${c}({x_{\tau}})>0$ for all $\tau\in\{0,1,\dots,t\}$. In other words, the RA measure ${g(\xi_{x_{0}}^{\pi,\phi},t)}$, defined as
${g(\xi_{x_{0}}^{\pi,\phi},t)}:=\min\Big{\{}r\big{(}{x_{t}}\big{)},\min_{\tau=0,\dots,t}c\big{(}{x_{\tau}}\big{)}\Big{\}},$
is positive, ${g(\xi_{x_{0}}^{\pi,\phi},t)}>0$, if and only if there exists a trajectory from $x_{0}$ reaching the target set safely.
An RA value function ${\bar{V}}(x)$ has been proposed in prior works [7, 8, 10, 9, 11, 15, 12], and it evaluates the maximum RA measure under the worst-case disturbance:
${\bar{V}}(x):=\max_{{\pi}}\min_{{\phi}}\sup_{t=0,\dots}{g(\xi_{x_{0}}^{\pi,\phi},t)}$ (8)
where ${\bar{V}}(x)>0$ if and only if $x\in{\mathcal{R}}$. We compute ${\bar{V}}(x)$ by solving its Bellman equation. However, ${\bar{V}}(x)$ has a non-contractive Bellman equation, whose solution may not recover ${\mathcal{R}}$, as shown in Figure 2. To address this, prior works include a time-discount factor $\gamma$ to create a contractive Bellman equation approximation. For each $\gamma\in(0,1)$, there is a unique solution to the approximated Bellman equation, which converges to ${\bar{V}}(x)$ as $\gamma$ is gradually annealed to 1.
Inspired by previous studies, we enhance computational efficiency by designing a new time-discounted RA value function. Ours incorporates a time-discount factor into the value function formulation, resulting in a contractive Bellman equation without the need for any approximation. This improvement eliminates the requirement of the $\gamma$-annealing process commonly used in prior works [7, 8, 10, 9, 11, 15, 12], and significantly reduces computation time.
The central part of our new value function is the time-discounted RA measure ${g_{\gamma}}({\xi_{x_{0}}^{\pi,\phi}},t)$, for a $\gamma\in(0,1)$,
${g_{\gamma}}({\xi_{x_{0}}^{\pi,\phi}},t):=\min\Big{\{}\gamma^{t}r\big{(}{x_{t}}\big{)},\min_{\tau=0,\dots,t}\gamma^{\tau}c\big{(}{x_{\tau}}\big{)}\Big{\}}.$
This yields a new time-discounted RA value function
${V_{\gamma}(x)}:=\max_{{\pi}}\min_{{\phi}}\sup_{t=0,\dots}{g_{\gamma}}({\xi_{x_{0}}^{\pi,\phi}},t).$ (9)
For all $\gamma\in(0,1)$ and any finite stage $t$, we have
${g_{\gamma}}({\xi_{x_{0}}^{\pi,\phi}},t)>0\Longleftrightarrow{g}({\xi_{x_{0}}^{\pi,\phi}},t)>0.$ (10)
Therefore, for all $\gamma\in(0,1)$, the super-zero level set of ${V_{\gamma}(x)}$, defined as ${\mathcal{V}_{\gamma}}:=\{x:{V_{\gamma}(x)}>0\}$, is equal to the RA set ${\mathcal{R}}$ in (4), and it includes all possible states that can reach the target set safely in finite time under the worst-case disturbance.
Figure 2: Comparing ${V_{\gamma}(x)}$ with ${\bar{V}}(x)$ from (8) and $V(x)$, a constructed solution to the Bellman equation of ${\bar{V}}(x)$ in prior works [7, 8, 10, 9, 11, 15, 12]. Consider a 1-dimensional dynamics: $x_{t+1}=1.01x_{t}+0.01(u_{t}+d_{t})$, with $|u_{t}|\leq 1$ and $|d_{t}|\leq 0.5$. We associate ${\mathcal{T}}=\{x:x<-1\}$ and ${\mathcal{C}}=\{x:x>-2\}$ with bounded, Lipschitz continuous functions $r(x)=\max(\min(-(x+1),10),-10)$ and $c(x)=\max(\min(x+2,10),-10)$, respectively. For all $\gamma\in(0,1)$, our super-zero level set $\{x:{V_{\gamma}(x)}>0\}$ equals the RA set =${\mathcal{R}}=\{x:-2<x<0.5\}$. By Theorem 2, ${V_{\gamma}(x)}$ is Lipschitz continuous if $\gamma\in(0,0.99009)$. The super-zero level set of ${\bar{V}}(x)$ also recovers ${\mathcal{R}}$, but ${\bar{V}}(x)$ is discontinuous at $x=0.5$ because the control fails to drive the state to ${\mathcal{T}}$ under the worst-case disturbance when $x_{t}\geq 0.5$. Finally, in the third subfigure, we show that the Bellman equation in prior works [7, 8, 10, 9, 11, 15, 12] has non-unique solutions, whose super-zero level set may not equal ${\mathcal{R}}$.
In what follows, we present the advantages of our new value function. First, we show that the Bellman equation for ${V_{\gamma}(x)}$ is a contraction mapping, with ${V_{\gamma}(x)}$ as its unique solution.
Theorem 1 (Contraction mapping).
Let $\gamma\in(0,1)$ and $V:\mathbb{R}^{n}\to\mathbb{R}$ be an arbitrary bounded function. Consider the Bellman operator ${B_{\gamma}}[\cdot]$ defined as,
${B_{\gamma}}[V(x)]\coloneqq\max_{u}\min_{d}\min\big{\{}c(x),\max\{r(x),\gamma V(f(x,u,d))\}\big{\}}.$ (11)
Then, ${V_{\gamma}(x)}$ in (9) is the unique solution to the Bellman equation $V(x)={B_{\gamma}}[V(x)]$. We have $\|{B_{\gamma}}[{V_{\gamma}^{1}}]-{B_{\gamma}}[{V_{\gamma}^{2}}]\|_{\infty}\leq\gamma\|{V_{\gamma}^{1}}-{V_{\gamma}^{2}}\|_{\infty}$, for all bounded functions ${V_{\gamma}^{1}}$ and ${V_{\gamma}^{2}}$.
Proof.
The proof can be found in the technical report [1]. ∎
Theorem 1 suggests that annealing $\gamma$ to 1 is unnecessary in our method because, for all $\gamma\in(0,1)$, our Bellman equation admits ${V_{\gamma}(x)}$ as the unique solution, and the super-zero level set of ${V_{\gamma}(x)}$ equals the ground truth RA set.
Furthermore, we show in the following result that our new value function can be constructed to be Lipschitz continuous, which facilitates efficient learning when approximating high-dimensional value functions using neural networks [40, 41].
Theorem 2 (Lipschitz continuity).
Suppose that $r(\cdot)$ and $c(\cdot)$ are $L_{r}$- and $L_{c}$-Lipschitz continuous functions, respectively. Assume also that the dynamics $f(x,u,d)$ is $L_{f}$-Lipschitz continuous in $x$, for all $u\in\mathcal{U}$ and $d\in\mathcal{D}$. Let $L:=\max(L_{r},L_{c})$. Then, ${V_{\gamma}(x)}$ is $L$-Lipschitz continuous if $\gamma L_{f}<1$.
Proof.
The proof can be found in the technical report [1]. ∎
The main idea of the proof is that a small perturbation in the state $x$ leads to a bounded change of the time-discounted RA measure value. Theorem 2 suggests that we can ensure the Lipschitz continuity of ${V_{\gamma}(x)}$ by selecting $\gamma<\frac{1}{{L_{f}}}$. In contrast, the classical RA value function ${\bar{V}}(x)$ can be discontinuous, as shown in Figure 2.
Moreover, the optimal control derived from our value function reaches the target set quickly, as a trajectory that reaches the target set rapidly incurs a high time-discounted RA measure value.
Theorem 3 (Fast reaching).
Let $x$ be in the RA set and ${\phi}$ be an arbitrary disturbance policy. Suppose $({\pi_{1}},{t_{1}})$ and $({\pi_{2}},{t_{2}})$ are two control policies and corresponding times to maximize the discounted RA measure $g_{\gamma}$, $\gamma\in(0,1]$,
$({\pi_{1}},{t_{1}}),({\pi_{2}},{t_{2}})\in\arg\max_{{\pi},t}g_{\gamma}(\xi_{x}^{{\pi},{\phi}},t).\vspace{-0.5em}$ (12)
If ${t_{1}}<{t_{2}}$, then for all $\check{\gamma}\in(0,\min\{\gamma,\frac{{V_{\gamma}(x)}}{\max_{x}r(x)}\})$, we have $g_{\check{\gamma}}(\xi_{x}^{{\pi_{2}},{\phi}},{t_{2}})<g_{\check{\gamma}}(\xi_{x}^{{\pi_{1}},{\phi}},{t_{1}})$.
Proof.
From the definitions of $\check{\gamma}$, ${g_{\gamma}}$, and ${V_{\gamma}(x)}$, and the boundedness of $r(x)$, we have $g_{\check{\gamma}}(\xi_{x}^{{\pi_{1}},{\phi}},{t_{2}})\leq\check{\gamma}^{{t_{1}}}\check{\gamma}\max_{x}r(x)\leq\check{\gamma}^{{t_{1}}}V_{\gamma}(x)<\big{(}\frac{\check{\gamma}}{\gamma}\big{)}^{{t_{1}}}V_{\gamma}(x)\leq\big{(}\frac{\check{\gamma}}{\gamma}\big{)}^{{t_{1}}}g_{\gamma}(\xi_{x}^{{\pi_{1}},{\phi}},{t_{1}})\leq g_{\check{\gamma}}(\xi_{x}^{{\pi_{1}},{\phi}},{t_{1}})$. This completes the proof. ∎
Theorem 3 also suggests that a control policy reaching the target set slowly may become suboptimal when $\gamma$ is decreased.
While a small time-discount factor in ${V_{\gamma}(x)}$ offers numerous benefits, it is not conclusive that $\gamma$ should always be near zero. In theory, for all $\gamma\in(0,1)$, the super-zero level set of ${V_{\gamma}(x)}$ recovers the exact RA set. However, in practice, a near-zero $\gamma$ can lead to a conservatively estimated RA set, where a trajectory reaching the target set at a late stage may have a near-zero or even negative time-discounted RA measure due to numerical errors. We will explore the trade-offs of selecting various $\gamma$ values in Section VII-D.
V Learning the New RA Value Function
Motivated by the fact that the optimal RA control policy is deterministic [36], we adapt max-min Deep Deterministic Policy Gradient (DDPG) [42], a deep RL method for learning deterministic policies and their value functions, to learn ${\pi}$, ${\phi}$ and ${V_{\gamma}(x)}$.
Let $\gamma$ be an arbitrary time discount factor in $(0,1)$. Similar to prior works [11, 15, 12], we approximate the optimal control policy $u={\pi}^{*}(x)$ and the worst-case disturbance policy $d={\phi}^{*}(x,{\pi}^{*}(x))$ by neural network (NN) policies $u={\pi_{\theta_{u}}}(x)$ and $d={\phi_{\theta_{d}}}(x)$, respectively, with ${\theta_{u}}$ and ${\theta_{d}}$ being their parameters. We define an NN Q function as ${Q_{\theta_{q}}}(x,u,d):\mathbb{R}^{n}\times{\mathcal{U}}\times{\mathcal{D}}\to\mathbb{R}$, where ${\theta_{q}}$ represents the NN’s parameter vector. Substituting ${\pi_{\theta_{u}}}$ and ${\phi_{\theta_{d}}}$ into ${Q_{\theta_{q}}}$, we can derive an NN value function ${V_{\theta}}(x):={Q_{\theta_{q}}}(x,{\pi_{\theta_{u}}}(x),{\phi_{\theta_{d}}}(x))$, where $\theta$ is the concatenation of parameters ${\theta_{q}}$, ${\theta_{u}}$ and ${\theta_{d}}$. Let $\mathbb{P}$ be the sampling distribution in learning. In max-min DDPG, we learn ${\pi_{\theta_{u}}}$, ${\phi_{\theta_{d}}}$ and ${Q_{\theta_{q}}}$ by jointly optimizing the following problems:
We learn ${\pi_{\theta_{u}}}$ by maximizing the Q value over ${\theta_{u}}$:
$\max_{{\theta_{u}}}\mathbb{E}_{x\sim{\mathbb{P}}}\ {Q_{\theta_{q}}}(x,{\pi_{\theta_{u}}}(x),{\phi_{\theta_{d}}}(x)).\vspace{-0.5em}$ (13)
We learn ${\phi_{\theta_{d}}}$ by minimizing the Q value over ${\theta_{d}}$:
$\min_{{\theta_{d}}}\mathbb{E}_{x\sim{\mathbb{P}}}\ {Q_{\theta_{q}}}(x,{\pi_{\theta_{u}}}(x),{\phi_{\theta_{d}}}(x)).\vspace{-0.5em}$ (14)
We learn ${Q_{\theta_{q}}}$ by minimizing the critic loss, also known as the Bellman equation error, over ${\theta_{q}}$:
$\min_{\theta_{q}}\mathbb{E}_{x\sim{\mathbb{P}}}\|{V_{\theta}}(x)-{B_{\gamma}}[{V_{\theta}}(x)]\|_{2}^{2}.$ (15)
We define the learned RA set $^{\hat{\mathcal{R}}}$ as the super zero-level set of ${V_{\theta}}(x)$. When DDPG converges to an optimal solution, ${V_{\theta}}(x)$ converges to ${V_{\gamma}(x)}$ due to Theorem 1. However, in practice, like other deep RL methods, DDPG often converges to a suboptimal solution with a near-zero critic loss. When ${V_{\theta}}(x)$ is a suboptimal solution, $^{\hat{\mathcal{R}}}$ cannot be reliably considered as the ground truth RA set ${\mathcal{R}}$. This motivates us to use a suboptimal learning result to certify a trustworthy RA set, as detailed in the following section.
VI Certifying RA Sets with Guarantees
In this section, we propose two methods to certify if a set of states belongs to the ground truth RA set. Both methods use a learned control policy, which is not necessarily optimal.
VI-A Certification using Lipschitz constants
We leverage a learned control policy ${\pi_{\theta_{u}}}$ and the Lipschitz constants of dynamics, reward and constraint to construct a theoretical lower bound of the ground truth value function ${V_{\gamma}}(x)$. If such a lower bound of ${V_{\gamma}}(x)$ is greater than zero for all states in the neighboring set of $x_{0}$, ${\mathcal{E}}_{x_{0}}:=\{x:\|x-x_{0}\|_{2}\leq\varepsilon_{x}\}$, then $V_{\gamma}(x)>0$, $\forall x\in{\mathcal{E}}_{x_{0}}$. We claim that the set ${\mathcal{E}}_{x_{0}}$ is within the ground truth RA set ${\mathcal{R}}$.
We begin constructing a lower bound of ${V_{\gamma}(x)}$ by considering a $T$-stage, disturbance-free, nominal trajectory $\{\bar{x}_{t}\}_{t=0}^{T}$,
${\bar{x}}_{t+1}=f({\bar{x}}_{t},{\bar{u}}_{t},0),{\bar{u}}_{t}={\pi_{\theta_{u}}}({\bar{x}}_{t}),\ \ \forall t=0,\dots,T-1$ (16)
and a state trajectory under ${\bar{u}}_{t}$ using an arbitrary $d_{t}\in{\mathcal{D}}$:
${\tilde{x}}_{t+1}=f({\tilde{x}}_{t},{\bar{u}}_{t},d_{t}),\ \ \forall t=0,\dots,T-1.$ (17)
Note that if we can verify that a trajectory starting at state $x$ reaches the target set safely despite any disturbance within $T$ stages, then it suffices to claim $x\in{\mathcal{R}}$. This certification horizon $T$ can be set arbitrarily. Ideally, we would set $T=\infty$, but it is impractical to evaluate an infinitely long trajectory. Therefore, during certification, we consider a finite, user-defined $T$. This $T$ should preferably be long enough to allow initial states to reach the target set. A short $T$ results in a conservative certification since it overlooks the possibility that the trajectory might safely reach the target set at a later time.
We assume that the dynamics $f(x,u,d)$ is Lipschitz continuous and there exists an upper bound on the disturbance in ${\mathcal{D}}$ at each stage $t$, i.e., $\|d_{t}\|_{2}\leq{\varepsilon_{d}}$. Let ${L_{f_{x}}}$ and ${L_{f_{d}}}$ be the Lipschitz constants of the dynamics $f$ with respect to the state $x$ and disturbance $d$, respectively. At time $t=1$, we observe
$\|{\bar{x}}_{1}$ $-{\tilde{x}}_{1}\|_{2}\leq\|f({\bar{x}}_{0},{\bar{u}}_{0},0)-f({\tilde{x}}_{0},{\bar{u}}_{0},d_{0})\|_{2}$
$\leq{L_{f_{x}}}\|{\bar{x}}_{0}-{\tilde{x}}_{0}\|_{2}+{L_{f_{d}}}\|0-d_{0}\|_{2}\leq{L_{f_{x}}}{\varepsilon_{x}}+{L_{f_{d}}}{\varepsilon_{d}}.$
At time $t=2$,
$\|{\bar{x}}_{2}-{\tilde{x}}_{2}\|_{2}$ $\leq\|f({\bar{x}}_{1},{\bar{u}}_{1},0)-f({\tilde{x}}_{1},{\bar{u}}_{1},d_{1})\|_{2}$
$\leq{L_{f_{x}}}\|{\bar{x}}_{1}-{\tilde{x}}_{1}\|_{2}+{L_{f_{d}}}{\varepsilon_{d}}.$
By induction, we have
$\|{\bar{x}}_{t}-{\tilde{x}}_{t}\|_{2}\leq{L_{f_{x}}^{t}}{\varepsilon_{x}}+\Sigma_{\tau=0}^{t-1}\ {L_{f_{x}}^{\tau}}{L_{f_{d}}}{\varepsilon_{d}}\ \ =:\Delta x_{t}.$ (18)
We define a convex outer approximation of the set of dynamically feasible states as ${\mathcal{X}_{t,\bar{x}_{0}}^{L}}:=\{x_{t}:\|x_{t}-{\bar{x}}_{t}\|_{2}\leq\Delta x_{t}\}$, and we check if for all $x_{t}\in{\mathcal{X}_{t,\bar{x}_{0}}^{L}}$, $r(x_{t})>0$ and $c(x_{t})>0$. By Lipschitz continuity of the reward function, we have,
$\forall x_{t}\in{\mathcal{X}_{t,\bar{x}_{0}}^{L}},\ \ \ \ \|r(\bar{x}_{t})-r(x_{t})\|_{2}\leq L_{r}\Delta x_{t}$
which yields a lower bound of $r(x_{t})$, for all $x_{t}\in{\mathcal{X}_{t,\bar{x}_{0}}^{L}}$:
${\check{r}_{t}^{L}}:=r(\bar{x}_{t})-L_{r}\Delta x_{t}\leq r(x_{t}).$ (19)
Similarly, we have a lower bound of $c(x_{t})$, for all $x_{t}\in{\mathcal{X}_{t,\bar{x}_{0}}^{L}}$:
${\check{c}_{t}^{L}}:=c(\bar{x}_{t})-L_{c}\Delta x_{t}\leq c(x_{t}).$ (20)
Using ${\check{r}_{t}^{L}}$ and ${\check{c}_{t}^{L}}$, we can construct a lower bound ${\check{V}^{L}_{\gamma}}({\bar{x}}_{0},T)$ for ${V_{\gamma}}(x_{0})$, for all $x_{0}\in{\mathcal{E}}_{{\bar{x}}_{0}}$:
${\check{V}^{L}_{\gamma}}({\bar{x}}_{0},T)$ $:=\max_{t=0,\dots,T}\min\{\gamma^{t}{\check{r}_{t}^{L}},\min_{\tau=0,\dots,t}\gamma^{\tau}{\check{c}_{\tau}^{L}}\}\leq{V_{\gamma}}(x_{0}).$ (21)
This implies
${\check{V}^{L}_{\gamma}}({\bar{x}}_{0},T)>0\Longrightarrow$ ${V_{\gamma}}(x_{0})>0,\forall x_{0}\in\mathcal{E}_{{\bar{x}}_{0}}.$ (22)
Thus, when ${\check{V}^{L}_{\gamma}}({\bar{x}}_{0},T)>0$, the set ${\mathcal{E}}_{{\bar{x}}_{0}}$ is certified to be within the ground truth RA set ${\mathcal{R}}$. Moreover, $\{{\bar{u}}_{t}\}_{t=0}^{T-1}$ are the certified control inputs that can drive all $x\in{\mathcal{E}}_{{\bar{x}}_{0}}$ to the target set ${\mathcal{T}}$ safely despite any disturbance in ${\mathcal{D}}$.
VI-B Certification using second-order cone programming
Lipschitz certification is fast to compute. However, the lower bound ${\check{V}^{L}_{\gamma}}$ can be conservative. In this subsection, we propose another certification method using second-order cone programming (SOCP), aiming to provide a less conservative RA certification. Our key idea is to construct SOCPs that search over a tight, convex outer approximation set of the dynamically feasible trajectories and verify whether the state trajectory reaches the target set safely under all disturbances, within a user-defined finite certification horizon $T$.
The construction of these SOCPs involves two steps.
First, we formulate a surrogate RA problem: A subset of the original target set ${\mathcal{T}}$, represented by the interior of a polytope ${\check{\mathcal{T}}}:=\{x:P_{i}x-k_{i}>0,i\in\mathcal{I}_{{\check{\mathcal{T}}}}\}$, is defined as a surrogate target set ${\check{\mathcal{T}}}$, where $\mathcal{I}_{\check{\mathcal{T}}}$ is a finite index set of the polytope’s edges. Additionally, we define a subset of the original constraint set ${\mathcal{C}}$, represented by the intersection of a finite number of quadratic functions’ super-zero level sets ${\check{\mathcal{C}}}:=\{x:\frac{1}{2}x^{\top}Q_{i}x+q_{i}^{\top}x+b_{i}>0,Q_{i}\succeq 0,i\in\mathcal{I}_{{\check{\mathcal{C}}}}\}$, as a surrogate (nonconvex) constraint set ${\check{\mathcal{C}}}$. The set ${\check{\mathcal{C}}}$ could be nonconvex when approximating collision avoidance constraints.
Subsequently, we leverage SOCPs to verify if a state trajectory from $x_{0}$ can reach ${\check{\mathcal{T}}}$ while staying within ${\check{\mathcal{C}}}$ despite any disturbance. We achieve this by sequentially minimizing each function defined in ${\check{\mathcal{T}}}$ and ${\check{\mathcal{C}}}$, iterating from the stage $\tau=0$ to $\tau=T$. If their minimum is positive at a stage $t$ and there is no intermediate stage $\tau<t$ such that $x_{\tau}$ is outside ${\check{\mathcal{C}}}$, we claim that the initial state $x_{0}$ can safely reach the original target set ${\mathcal{T}}$, under all possible disturbances.
To be more specific, we can check if there exists a stage $t\in\{0,1,\dots,T\}$ such that, for all disturbances, all dynamically feasible states $x_{t}$, originating from the initial states set $\mathcal{E}_{{\bar{x}}_{0}}:=\{x:\|x-{\bar{x}}_{0}\|_{2}\leq{\varepsilon_{x}}\}$, are within ${\check{\mathcal{T}}}$ and for all stages $\tau\leq t$, $x_{\tau}$ are within ${\check{\mathcal{C}}}$. If this condition is met, we claim that $\mathcal{E}_{{\bar{x}}_{0}}$ is within the ground truth RA set ${\mathcal{R}}$, i.e., $\mathcal{E}_{{\bar{x}}_{0}}\subseteq\mathcal{R}$. Otherwise, some of its elements could be outside ${\mathcal{R}}$, and therefore we do not claim $\mathcal{E}_{{\bar{x}}_{0}}\subseteq{\mathcal{R}}$. To make the analysis tractable, we define the nominal state trajectory $\{{\bar{x}}_{t}\}_{t=0}^{T}$ and nominal control trajectory $\{{\bar{u}}_{t}\}_{t=0}^{T-1}$ as in (16). The nominal state and control trajectories allow us to formulate a convex set ${\mathcal{X}_{t,{\bar{x}}_{0}}^{S}}$ for outer approximating the set of dynamically feasible states:
${\mathcal{X}_{t,{\bar{x}}_{0}}^{S}}:=\{x_{t}:\exists\{d_{\tau}\}_{\tau=0}^{t-1}\textrm{ and }x_{0}\textrm{ such that }\forall\tau\leq t-1,$
$x_{\tau+1}\leq{\hat{A}}_{\tau}x_{\tau}+{\hat{B}}_{\tau}{\bar{u}}_{\tau}+{\hat{D}}_{\tau}d_{\tau}+{\hat{c}}_{\tau},{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(\textrm{Upper bound on $f$})}$
$x_{\tau+1}\geq{\check{A}}_{\tau}x_{\tau}+{\check{B}}_{\tau}{\bar{u}}_{\tau}+{\check{D}}_{\tau}d_{\tau}+{\check{c}}_{\tau},{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(\textrm{Lower bound on $f$})}$
$\|d_{\tau}\|_{2}\leq\varepsilon_{d},\hskip 103.00018pt{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(\textrm{Disturbance bound})}$
$\|x_{0}-{\bar{x}}_{0}\|_{2}\leq\varepsilon_{x}\}\hskip 80.00012pt{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(\textrm{Initial statebound})}$
where the bounds on the dynamics $f$ can be derived using a Taylor series. At a stage $t$, we can verify if $x_{t}\in{\check{\mathcal{T}}}$ by solving a sequence of SOCPs iterating over all $i\in\mathcal{I}_{\check{\mathcal{T}}}$, and checking if their minimum is positive:
${\check{r}_{t}^{S}}:=\min_{i\in\mathcal{I}_{{\check{\mathcal{T}}}}}\big{\{}\min_{x_{t}\in{\mathcal{X}_{t,{\bar{x}}_{0}}^{S}}}$ $P_{i}^{\top}x_{t}-k_{i}\big{\}}.$ (23)
Similarly, for checking whether $x_{t}\in{\check{\mathcal{C}}}$, we can evaluate if the following term is positive:
${\check{c}_{t}^{S}}:=\min_{i\in\mathcal{I}_{{\check{\mathcal{C}}}}}\big{\{}\min_{x_{t}\in{\mathcal{X}_{t,{\bar{x}}_{0}}^{S}}}\frac{1}{2}x_{t}^{\top}Q_{i}x_{t}+q_{i}^{\top}x_{t}+b_{i}\big{\}}.$ (24)
Combining the above two terms, we can construct a conservative certificate ${\check{V}^{S}_{\gamma}}$ for verifying if a state $\bar{x}_{0}$ and its neighboring set $\mathcal{E}_{\bar{x}_{0}}$ are within the ground truth RA set ${\mathcal{R}}$,
${\check{V}^{S}_{\gamma}}(\bar{x}_{0},T):=\max_{t=0,\dots,T}\min\{\gamma^{t}{\check{r}_{t}^{S}},\min_{\tau=0,\dots,t}\gamma^{\tau}{\check{c}_{\tau}^{S}}\}.$ (25)
This suggests
${\check{V}^{S}_{\gamma}}({\bar{x}}_{0},T)>0\Longrightarrow{V_{\gamma}}(x_{0})>0,\forall x_{0}\in\mathcal{E}_{{\bar{x}}_{0}}$ (26)
and $\{{\bar{u}}_{t}\}_{t=0}^{T-1}$ are the certified control inputs, capable of driving all $x\in{\mathcal{E}}_{{\bar{x}}_{0}}$ to the target set ${\mathcal{T}}$ safely despite any disturbance.
Running example (continued). We evaluate ${\check{r}_{t}^{S}}=\min_{i}\{\check{r}_{t,i}^{S}\}_{i=1}^{6}$ by solving the following SOCPs, where each $\check{r}_{t,i}^{S}$ corresponds a function in the definition of ${\mathcal{T}}$ in (5):
$\check{r}_{t,1}^{S}=\min_{x_{t}\in{\mathcal{X}_{t,{\bar{x}}_{0}}^{S}}}p_{y,t}^{1}-p_{y,t}^{2},$ $\check{r}_{t,2}^{S}=\min_{x_{t}\in{\mathcal{X}_{t,{\bar{x}}_{0}}^{S}}}v_{y,t}^{1}-v_{y,t}^{2}$
$\check{r}_{t,3}^{S}=\min_{x_{t}\in{\mathcal{X}_{t,{\bar{x}}_{0}}^{S}}}0.3-p_{x,t}^{1},$ $\check{r}_{t,4}^{S}=\min_{x_{t}\in{\mathcal{X}_{t,{\bar{x}}_{0}}^{S}}}0.3+p_{x,t}^{1}$
$\check{r}_{t,5}^{S}=\min_{x_{t}\in{\mathcal{X}_{t,{\bar{x}}_{0}}^{S}}}0.3-p_{y,t}^{1},$ $\check{r}_{t,6}^{S}=\min_{x_{t}\in{\mathcal{X}_{t,{\bar{x}}_{0}}^{S}}}0.3+p_{y,t}^{1}$
Similarly, we can evaluate ${\check{c}_{t}^{S}}=\min_{i}\{\check{c}_{t,i}^{S}\}_{i=1}^{5}$ by considering the following SOCPs, where each $\check{c}_{t,i}^{S}$ corresponds to a constraint function in (6) and (7):
$\check{c}_{t,i}^{S}=\min_{x_{t}\in{\mathcal{X}_{t,{\bar{x}}_{0}}^{S}}}(-1)^{i}\times p_{x,t}^{1}-p_{y,t}^{1}+0.05,\hskip 5.69046pti\in\{1,2\},$
$\check{c}_{t,i}^{S}=\min_{x_{t}\in{\mathcal{X}_{t,{\bar{x}}_{0}}^{S}}}(-1)^{i}\times p_{z,t}^{1}-p_{y,t}^{1}+0.05,\hskip 5.69046pti\in\{3,4\},$
We overapproximate the maximum height difference between two drones at the stage $t$ via $\Delta_{z,t}^{21}:=\max_{x_{t}\in{\mathcal{X}_{t,{\bar{x}}_{0}}^{S}}}p_{z,t}^{2}-p_{z,t}^{1}$, and consider
$\check{c}_{t,5}^{S}=\min_{x_{t}\in{\mathcal{X}_{t,{\bar{x}}_{0}}^{S}}}\mathopen{}\mathclose{{}\left\|\begin{bmatrix}p_{x,t}^{1}-p_{x,t}^{2}\\ p_{y,t}^{1}-p_{y,t}^{2}\end{bmatrix}}\right\|_{2}^{2}-(1+\max(\Delta_{z,t}^{21},0))\times 0.2$ (27)
Remark 1 (Comparing the two methods).
1) SOCP certification can be less conservative than Lipschitz certification when ${\check{\mathcal{T}}}$ and ${\check{\mathcal{C}}}$ can represent ${\mathcal{T}}$ and ${\mathcal{C}}$ exactly, and ${\mathcal{X}_{t,{\bar{x}}_{0}}^{S}}$ is a subset of the Lipschitz dynamically feasible set ${\mathcal{X}_{t,\bar{x}_{0}}^{L}}$, as defined in Section VI-A, for all $t\leq T$. This is because Lipschitz certification adds extra conservatism when estimating lower bounds of $r(x)$ and $c(x)$ using the Lipschitz constant, as shown in (19) and (20); 2) However, Lipschitz certification is faster to compute than SOCP because calculating ${\check{r}_{t}^{L}}$ and ${\check{c}_{t}^{L}}$ is easier than evaluating ${\check{r}_{t}^{S}}$ and ${\check{c}_{t}^{S}}$.
Remark 2.
The computational complexity of evaluating ${\check{V}^{L}_{\gamma}}(x,T)$ and ${\check{V}^{S}_{\gamma}}(x,T)$ scales polynomially with both the dimension of the dynamical system and the length of $T$.
VI-C Combining reachability learning and certification for computing trustworthy RA sets
We integrate the reachability learning and certification into a new framework of computing trustworthy RA sets, as described in Algorithm 1. The super zero-level set of ${V_{\theta}}$ provides an estimation of the ground truth RA set. We use ${\pi_{\theta_{u}}}$ to certify a set of states, ensuring deterministic RA guarantees there. In particular, we can apply certification either online or offline:
Online certification: Let $x$ be an arbitrary state. We can use the RA certificates ${\check{V}^{L}_{\gamma}}(x,T)$ in (21) or ${\check{V}^{S}_{\gamma}}(x,T)$ in (25) as online RA certification methods, verifying if all elements in ${\mathcal{E}}_{x}=\{x^{\prime}:\|x^{\prime}-x\|_{2}\leq{\varepsilon_{x}}\}$ can reach the target set safely despite any disturbance. We can compute ${\check{V}^{L}_{\gamma}}(x,T)$ and ${\check{V}^{S}_{\gamma}}(x,T)$ in real-time (10 Hz or faster), as shown in Figure 5. It can also be integrated into the safety filter proposed in [11], and our method provides a more robust RA certification than [11] as we certify whether all elements in ${\mathcal{E}}_{x}$ can reach the target set safely, allowing us to verify the RA capability without perfect state estimation.
Offline certification: We consider a finite set of states :=$\mathcal{L}:=\{x^{(i)}\}_{i=1}^{N}$ such that the union of their neighboring sets ${\mathcal{E}}_{x^{(i)}}=\{x:\|x-x^{(i)}\|_{2}\leq{\varepsilon_{x}}\}$ covers the set of states that we aim at certifying. For example, this includes the area near the orange gate in drone racing, as show in Figure 1. We enumerate each element $x\in\mathcal{L}$ and certify whether $\mathcal{E}_{x}\subseteq{\mathcal{R}}$ by checking if ${\check{V}^{L}_{\gamma}}$ in (21) or ${\check{V}^{S}_{\gamma}}$ in (25) is positive. The union ${\mathcal{S}}$ of those certified sets constitutes a subset of the ground truth RA set ${\mathcal{R}}$. For all elements in ${\mathcal{S}}$, we guarantee that they can reach the target set safely amid potential disturbances.
Require:  an arbitrary $\gamma\in(0,1)$, a finite list of states $\mathcal{L}$, and a certification horizon $T<\infty$;
Initialization: certified RA set ${\mathcal{S}}\leftarrow\{\}$;
Learn ${\pi_{\theta_{u}}}$, ${\phi_{\theta_{d}}}$ and ${V_{\theta}}$ via max-min DDPG [42];
// Certifications:
for $x_{0}\in\mathcal{L}$ do
       if ${\check{V}^{L}_{\gamma}}(x_{0},T)>0$ or ${\check{V}^{S}_{\gamma}}(x_{0},T)>0$ then
            ${\mathcal{S}}\leftarrow{\mathcal{S}}\cup\{x:\|x_{0}-x\|_{2}\leq{\varepsilon_{x}}\}$
return  certified RA set ${\mathcal{S}}$
Algorithm 1 Reachability Learning and Certifications:
VII Experiments
We test our reachability learning and certification methods [^1] Experiment code and details are available at https://github.com/jamesjingqili
/Lipschitz_Continuous_Reachability_Learning.git
in a 12-dimensional drone racing hardware experiment, and a triple-vehicle highway take-over simulation. In the highway simulation, we control one ego vehicle, modeled with nonlinear unicycle dynamics, to safely overtake another vehicle while avoiding a third vehicle driving in the opposite direction, as shown in Figure 4. Figures 1 and 4 demonstrate the high quality of the learned and certified RA sets. Additionally, the hardware experiments in Figure 3 show that our learned control policy ${\pi_{\theta_{u}}}$ safely controls the ego drone (highlighted with brightness) to overtake the other drone when the initial states of both agents are within the SOCP certified set.
VII-A Hypothesis 1: Our learned policy has a higher success rate than state-of-the-art constrained RL methods
We compare our learned policy ${\pi_{\theta_{u}}}$ with Deep Deterministic Policy Gradient-Lagrangian (DDPG-L)[34], Soft Actor Critic-Lagrangian (SAC-L) [35], Constrained Policy Optimization (CPO) [33]. We summarize the results in Table I. The success rate is estimated by computing the ratio of sampled initial states that can reach the target set safely under randomly generated disturbances. Our method achieves a 1.0 success rate when the initial states are sampled from the SOCP certified set, validating the deterministic guarantee that all elements in the certified sets can safely reach the target set despite any disturbance. However, CPO fails to converge for the drone racing experiment due to the complex and nonconvex constraints.
TABLE I: Success rates table. Our method achieves a $1.0$ success rate when the initial states are sampled from the SOCP certified set.
Figure 3: In both subfigures, the initial states are within the SOCP certified set shown in Figure 1, validating the deterministic guarantees that the ego drone can safely overtake the other despite any disturbance on the other drone’s acceleration, with $\|d_{t}\|_{2}\leq 0.1m/s^{2}$. The remaining 9-dimensional initial state includes $[v_{x,t}^{1},v_{y,t}^{1},v_{z,t}^{1},p_{x,t}^{2},p_{y,t}^{2},p_{z,t}^{2},v_{x,t}^{2},v_{y,t}^{2},v_{z,t}^{2}]=[0,0.7,0,0.4,-2.2,0,0,0.3,0]$.
Figure 4: Highway reachability analysis: In (a), we simulate the nonlinear dynamics with the learned policy ${\pi_{\theta_{u}}}$ and randomly sampled disturbances on other vehicles’ acceleration. The 10-dimensional state space includes $[p_{x,t}^{1},p_{y,t}^{1},v_{t}^{1},\theta_{t}^{1},p_{x,t}^{2},p_{y,t}^{2},v_{y,t}^{2},p_{x,t}^{3},p_{y,t}^{3},v_{y,t}^{3}]$. In (b), we project our learned value function, with $\gamma=0.95$, onto the $(x,y)$ position of the ego vehicle. In (c), we plot the RA set learned using the state-of-the-art method [12, 15] with $\gamma=0.95$. As suggested in [8], annealing $\gamma\to 1$ is necessary for prior works; otherwise, the learned RA sets in prior works are conservative. In (d), we plot our certified RA sets.
Figure 5: Histogram of the time required for computing ${\check{V}^{L}_{\gamma}}(x,T)$ and ${\check{V}^{S}_{\gamma}}(x,T)$ for each of the 10,000 randomly sampled states $x$. The certification horizons for drone racing and highway are $T=15$ and $T=30$, respectively.
VII-B Hypothesis 2: Our online RA set certification methods can be computed in real-time
Figure 5 shows that ${\check{V}^{L}_{\gamma}}(x,T)$ and ${\check{V}^{S}_{\gamma}}(x,T)$ can be computed in real-time to certify if all elements in ${\mathcal{E}}_{x}=\{x^{\prime}:\|x^{\prime}-x\|_{2}\leq 0.1\}$ can safely reach the target set, amid all potential disturbances. This enables real-time online certification.
Figure 6: The volumes of the learned RA set, SOCP certified set, and the Lipschitz certified set change as $\gamma$ varies. We estimate the set volumes using the Monte Carlo method with 10,000 random samples in the state space.
Figure 7: The average time taken for reaching the target set grows as $\gamma$ increasing.
Figure 8: The convergence of the critic loss of max-min DDPG under our new Bellman equation (Ours) and under the approximated Bellman equation (Baseline) from prior works [15, 12], with different $\gamma$ values. Our critic loss under $\gamma=0.95$ converges more rapidly than under $\gamma=0.9999$ possibly due to the Lipschitz continuity of ${V_{\gamma}(x)}$ when $\gamma=0.95$. Both the baseline and ours are trained using the same training parameters and random initializations. For drone racing, we parameterize each of ${\pi_{\theta_{u}}}$, ${\phi_{\theta_{d}}}$ and ${Q_{\theta_{q}}}$ using a 4-layer ReLU NN, with 512 neurons in each layer. For the highway example, we parameterize each ${\pi_{\theta_{u}}}$, ${\phi_{\theta_{d}}}$ and ${Q_{\theta_{q}}}$ using a 3-layer ReLU NN, with 512 neurons in each layer. The NN training is conducted under 5 different random seeds.
Figure 9: The Lipschitz continuity of our new RA value functions appears to accelerate learning, and this phenomenon is not limited to max-min DDPG. In the first subfigure, we apply max-min DDPG under our new Bellman equation (Ours) and the approximated Bellman equation (Baseline) from prior works [12, 15]. In the second subfigure, we apply max-min SAC [12, 15] under both Bellman equations. The dynamics considered are the same as in Figure 2, with $r(x)=\max(\min(2-x^{2},10),-10)$ and $c(x)=\max(\min(x+1.5,10),-10)$. Due to the simplicity of this problem, we model ${Q_{\theta_{q}}}$ as a 2-layer ReLU NN with 16 neurons in each layer, and model both ${\pi_{\theta_{u}}}$ and ${\phi_{\theta_{d}}}$ as 2-layer ReLU NNs, each with 4 neurons per layer. We train the NNs under 5 different random seeds. Under $\gamma=0.95$, ${V_{\gamma}(x)}$ is Lipschitz continuous and our method converges rapidly. However, as $\gamma\to 1$, our learning process becomes unstable possibly due to the loss of Lipschitz continuity of ${V_{\gamma}(x)}$.
VII-C Hypothesis 3: The Lipschitz continuity of our new value function appears to accelerate learning
In Figures 8 and 9, we compare the critic loss (defined in (15) to measure the Bellman equation error) under different Bellman equations and time-discount factor $\gamma$ values. We observe that the critic loss under our new Bellman equation converges rapidly when $\gamma$ is chosen to ensure the Lispchitz continuity of our new value function.
VII-D The trade-off of selecting a time-discount factor $\gamma$
We summarize our result in Figures 6 and 7. With a small $\gamma$, the learned RA sets can be conservative due to numerical errors, as an initial state whose optimal trajectory reaches the target set at a later stage may have near-zero or even negative time-discounted RA measures. However, the optimal policy tends to drive the state to the target set rapidly, as depicted in Figure 7. Conversely, when $\gamma$ is high, the value function could be discontinuous, leading to unstable learning and resulting in a suboptimal value function. In the drone racing and highway experiments, we find that $\gamma=0.95$ ensures the Lipschitz continuity of ${V_{\gamma}(x)}$, thereby enhancing learning efficiency, and also mitigates unnecessary conservatism.
VIII Conclusion
We propose a new framework for learning trustworthy RA sets. Our method features a newly designed time-discounted RA value function, which is provably Lipschitz continuous and has a Bellman operator that is a contraction mapping. This enhances computational efficiency by eliminating the need for the time-discount factor annealing process used in previous works. We use max-min DDPG to learn our value functions and propose two efficient methods to certify if a set of states can safely reach the target set despite any disturbance, with deterministic guarantees. We validate our methods in drone racing hardware experiments and highway take-over simulations. In future research, we plan to extend this work to online reachability learning in uncertain environments and propose more efficient and less conservative reachability set certification methods by leveraging the problem structure.
IX Acknowledgement
We appreciate Andrea Bajcsy and David Fridovich-Keil for valuable discussion.